# Bridging Swarm Intelligence and Reinforcement Learning

### v1

@@ Abstract.
Swarm intelligence (SI) explores how large groups of simple individuals (e.g., insects, fish, birds) collaborate to produce complex behaviors, exemplifying that the whole is greater than the sum of its parts. A fundamental task in SI is Collective Decision-Making (CDM), where a group selects the best option among several alternatives, such as choosing an optimal foraging site. In this work, we demonstrate a theoretical and empirical equivalence between CDM and single-agent reinforcement learning (RL) in multi-armed bandit problems, utilizing concepts from opinion dynamics, evolutionary game theory, and RL. This equivalence bridges the gap between SI and RL and leads us to introduce a novel abstract RL update rule called Maynard-Cross Learning. Additionally, it provides a new population-based perspective on common RL practices like learning rate adjustment and batching. Our findings enable cross-disciplinary fertilization between RL and SI, allowing techniques from one field to enhance the understanding and methodologies of the other.
Key words and phrases:
Swarm Intelligence, Reinforcement Learning, Evolutionary Game Theory, Opinion Dynamics
@@ 1. Introduction
In this paper, we focus the specific SI problem of Collective Decision Making (CDM). In CDM, individuals work together to reach an agreement on the best option from a set of alternatives, a problem commonly called the best-of-n decision problem. To solve this problem, researchers have turned to opinion dynamics Xia et al. (2011), a field that studies how opinions spread in a population. In particular, in the voter rule Castellano et al. (2009); Redner (2019), an individual copies the opinion of a randomly chosen neighbor. Similarly, researchers have taken inspiration from the house-hunting behavior of honey bees to create the weighted voter rule Valentini et al. (2014); Reina et al. (2024). In this rule, after scouting one of $n$ potential nesting areas, bees come back to perform a “dance” Visscher and Camazine (1999) that describes the coordinates of the option that they have explored. According to the weighted voter model, this dance is performed at a frequency that is proportional to the estimated quality of the explored area. Other bees go scout the area corresponding to the first dance they witness, and this process repeats until the entire colony converges to the same option.
@@
Next, we turn toward reinforcement learning (RL), where an agent [^1] To avoid confusion, we use “agent” in the context of RL and “individual” in the context of a population, wherever possible learns to solve a task by interacting with the environment to maximize a reward signal Sutton and Barto (2018). RL has been successfully applied to solve complex problems in various fields such as robotics Kaufmann et al. (2023), nuclear fusion Seo et al. (2024), and games Mnih et al. (2013). In this paper, we are specifically interested in multi-armed bandits Sutton and Barto (2018), in which a single agent makes choices among different options (or “arms”) to maximize its reward. Among the many learning algorithms designed to solve this task (Upper-confidence-Bound Auer et al. (2002) (UCB), $\epsilon$-greedy Sutton and Barto (2018), Gradient Bandit Williams (1992)…), we consider the Cross Learning Cross (1973) update rule, closely related to the Gradient Bandit algorithm.
@@
Although SI and RL are seemingly disjoint, we show that these fields can in fact be bridged via the Replicator Dynamic Sandholm et al. (2008) (RD), a famous equation used in Evolutionary Game Theory (EGT) to model the outcome of evolutionary processes through the idea of survival of fittest. In the rest of the paper, we demonstrate a mathematical equivalence between different concepts from SI and RL:
@@
  • We first show that a large population whose members follow the voter rule can be seen as a single abstract RL agent following the Cross Learning update rule.
@@
  • Next, via a similar argument, we show that the weighted voter rule, yields another abstract RL update that we coin Maynard-Cross Learning.
@@ 2.2. Evolutionary game theory
Evolutionary game theory (EGT) studies population games Sandholm (2010). In a single-population game, a population $\mathcal{P}$ is made of a large number of individuals, where any individual $i$ is associated with a type, denoted by $T_{i}\in\{1,\dots,n\}$. The population vector $\pi=(\pi_{1},\dots,\pi_{n})$ represents the fraction of individuals in each type ($\sum_{i}\pi_{i}=1$). Individuals are repeatedly paired at random to play a game, each receiving a separate payoff defined by the game bi-matrix [^2] A bimatrix game is defined by a pair of identical-shape matrices, one per agent. $A$. Individuals adapt their type based on these payoffs according to an update rule. [^3] In population games, an update rule is sometimes called a revision protocol. One notable such rule is imitation of success Sandholm (2010):
@@
where $\dot{\pi}_{a}$ is the derivative of the $a$-th component of the population vector $\pi_{a}$, $q^{\pi}_{a}:=\mathbb{E}[r_{a}]$ is the expected payoff of the type $a$ against the current population, and $v^{\pi}:=\sum_{b}\pi_{b}\mathbb{E}[r_{b}]$ is the current average payoff of the entire population. [^6] In Evolutionary Game Theory, expected payoffs are often referred to as “fitness”, as they model the reproductive fitness of the different types. Further, we describe another useful variant of the TRD for later convenience in the paper, the Maynard Smith Replicator Dynamic Smith (1982) (MRD):
@@ 3.2. Learning rate and batch-size
The population-policy equivalence gives an interesting interpretation to the learning rate $\alpha$ commonly used in RL. Under the population perspective, $\alpha$ describes the number of individuals in the population. In Sec. 5, we empirically show that the CL update rule described in Eq. 1 does not typically converge to the optimal action. However, using a small enough learning rate (i.e., a large enough population size) alleviates this issue [^9] As implied by Eq. 14, assuming that, when performed sequentially on randomly sampled individuals instead of parallelly in one step across the entire population, the voter rule still yields the Replicator Dynamic, which we conjecture. We leave a proof of this conjecture for future work..
@@ 4. Methods
We present two types of experiments to validate the findings from the previous section. First, we implement the two RL update rules, CL and MCL in two variants: batched and non-batched. Second, we conduct population-based experiments using $R_{\text{voter}}$ and $R_{\text{wvoter}}$ (VR, WVR) for different population sizes. Moreover, we also numerically simulate the TRD and MRD to show how the above experiments compare with the analytical solutions.
@@ 4.2. RL experiments
Moreover, since this update rule can make $\pi$ invalid, i.e., components could become negative or above one (see footnote 11), we clamp $\pi$ between 0 and 1:
@@
We call this training step a run, and perform $R$ runs per seed.
Batched: In these experiments, we implement the batched variants of update rules CL (Eq. 2) and MCL (Eq. 25), henceforth named B-CL and B-MCL. B-CL is a straightforward batching of the CL update rule, averaging over a batch of $B$ samples instead of one sample to update $\pi$. Whereas, with B-MCL, $v^{\pi}$ is no longer a moving average of the rewards but rather the mean of batch rewards. We also need to explicitly clamp the policy between 0 and 1 to ensure it remains valid. B-MCL also uses $B$ samples simultaneously similar to B-CL. Similar to non-batched experiments, we perform $R$ runs per seed.
@@ 5. Results
Finally, from the discussions related to batched RL updates and population experiments, we empirically validate Proposition 1 and Proposition 2.
@@ 6. Conclusions
With Propositions 1 and 2, we have demonstrated how RD is the underlying connection between Reinforcement Learning and Collective Decision-Making. Further, we have empirically validated this correspondence. This correspondence opens a bridge between these two fields, enabling the flow of ideas, new perspectives, and analogies for both. For example, it can be seen that Cross Learning, Maynard-Cross Learning, and, more generally, Reinforcement Learning take a single-agent perspective, where information from consecutive action samples/batches is accumulated into one centralized agent’s policy. On the other hand, $R_{\text{voters}}$ and $R_{\text{wvoters}}$ take a multi-agent perspective, where every individual implements simple local and decentralized rules, making independent decisions, which leads to an emergent collective policy.
@@
Significance for SI. The population-policy equivalence highlights how certain Swarm Intelligence methods are equivalent to single-agent Reinforcement Learning, demonstrating agency of the entire swarm as a single learning entity. Therefore, one could imagine that Multi-Agent Reinforcement Learning would similarly yield equivalent multi-swarm systems, where two or more coexisting swarms would compete/collaborate for resources (i.e., prisoners dilemma, hawk dove, etc). Further, ideas that empower Reinforcement Learning, could be ported to swarm intelligence and swarm robotics.
@@ A.1. Batched RL experiments
Figure 5. Batched RL figures for large batch size and the two other environments where $q^{pi}_{a}$’s are near zero and near one.
@@ A.2. Population experiments
Figure 6. Population experiments for large population size and the two other environments where $q^{pi}_{a}$’s are near zero and near one.
@@ A.3. Hyperparametrs
final time ($t_{f}$) $\alpha\times R$ = 100

### v2

@@ Abstract.
Swarm intelligence (SI) explores how large groups of simple individuals (e.g., insects, fish, birds) collaborate to produce complex behaviors, exemplifying that the whole is greater than the sum of its parts. A fundamental task in SI is Collective Decision-Making (CDM), where a group selects the best option among several alternatives, such as choosing an optimal foraging site. In this work, we demonstrate a theoretical and empirical equivalence between CDM and single-agent reinforcement learning (RL) in multi-armed bandit problems, utilizing concepts from opinion dynamics, evolutionary game theory, and RL. This equivalence bridges the gap between SI and RL and leads us to introduce a novel biologically plausible RL update rule called Maynard-Cross Learning. Additionally, it provides a new population-based perspective on common RL practices like learning rate adjustment and batching. Our findings enable cross-disciplinary fertilization between RL and SI, allowing techniques from one field to enhance the understanding and methodologies of the other.
@@ 1. Introduction
In this paper, we focus the specific SI problem of Collective Decision Making (CDM). In CDM, individuals work together to reach an agreement on the best option from a set of alternatives, a problem commonly called the best-of-n decision problem. Due to its straightforward and generic framework, CDM has proven effective for modeling decision-making problems in diverse domains, such as honeybee colonies Reina et al. (2015); Bose et al. (2017), human societies Jackson and Golub (2010), and robot swarms Valentini et al. (2016, 2014). To solve this problem, researchers have turned to opinion dynamics Xia et al. (2011), a field that studies how opinions spread in a population. In particular, in the voter rule Castellano et al. (2009); Redner (2019), an individual copies the opinion of a randomly chosen neighbor. Similarly, researchers have taken inspiration from the house-hunting behavior of honey bees to create the weighted voter rule Valentini et al. (2014); Reina et al. (2024). In this rule, after scouting one of $n$ potential nesting areas, bees come back to perform a “dance” Visscher and Camazine (1999) that describes the coordinates of the option that they have explored. According to the weighted voter model, this dance is performed at a frequency that is proportional to the estimated quality of the explored area. Other bees go scout the area corresponding to the first dance they witness, and this process repeats until the entire colony converges to the same option. Further, investigations related to dynamic qualities for options Prasetyo et al. (2019), multi-variable qualities Ebert et al. (2018), continuous space options Raoufi et al. (2021), Bayesian approaches to model beliefs Ebert et al. (2020), and quality magnitude sensitivity Pirrone et al. (2022) have been carried out in the literature.
@@
Next, we turn toward reinforcement learning (RL), where an agent [^1] To avoid confusion, we use “agent” in the context of RL and “individual” in the context of a population, wherever possible learns to solve a task by interacting with the environment to maximize a reward signal Sutton and Barto (2018). RL has been successfully applied to solve complex problems in various fields such as robotics Kaufmann et al. (2023), nuclear fusion Seo et al. (2024), and games Mnih et al. (2013). In this paper, we are specifically interested in multi-armed bandits Sutton and Barto (2018), in which a single agent makes choices among different options (or “arms”) to maximize its reward. Among the many learning algorithms designed to solve this task (Upper-confidence-Bound Auer et al. (2002) (UCB), $\epsilon$-greedy Sutton and Barto (2018), Gradient Bandit Williams (1992), etc.), we consider the Cross Learning Cross (1973) update rule, closely related to the Gradient Bandit algorithm. The Gradient Bandit algorithm introduces the foundational concept of policy gradient optimization, which forms the basis for advanced policy-based RL algorithms like SAC, PPO, and REINFORCE, tailored for sequential decision-making tasks in complex environments.
@@
Although SI and RL are seemingly disjoint, we show that these fields can in fact be bridged via the Replicator Dynamic Sandholm et al. (2008) (RD), a famous equation used in Evolutionary Game Theory (EGT) to model the outcome of evolutionary processes through the idea of survival of fittest. While previous works have explored connections between decision-making in honeybee and distributed human brain cognition Passino et al. (2008) (hive mind), to the best of our knowledge, our work is the first to establish parallels between decision-making in honeybees and reinforcement learning. In the rest of the paper, we demonstrate the mathematical equivalence between different concepts from SI and RL:
@@
  • We first show that a large non-learning population whose members follow the voter rule can be seen as a single abstract RL agent following the Cross Learning update rule.
@@
  • Next, via a similar argument, we show that the weighted voter rule, yields a novel biologically plausible RL update rule that we coin Maynard-Cross Learning.
@@ 2.2. Evolutionary game theory
Evolutionary game theory (EGT) studies population games Sandholm (2010). In a single-population game, a population $\mathcal{P}$ is made of a large number of individuals, where any individual $i$ is associated with a type, denoted by $T_{i}\in\{1,\dots,n\}$. The population vector $\pi=(\pi_{1},\dots,\pi_{n})$ represents the fraction of individuals in each type ($\sum_{i}\pi_{i}=1$). Individuals are repeatedly paired at random to play a game, each receiving a separate payoff defined by the game bi-matrix $A$. Individuals adapt their type based on these payoffs according to an update rule. One notable such rule is imitation of success Sandholm (2010):
@@
where $\dot{\pi}_{a}$ is the derivative of the $a$-th component of the population vector $\pi_{a}$, $q^{\pi}_{a}:=\mathbb{E}[r_{a}]$ is the expected payoff of the type $a$ against the current population, and $v^{\pi}:=\sum_{b}\pi_{b}\mathbb{E}[r_{b}]$ is the current average payoff of the entire population. Further, we describe another useful variant of the TRD for later convenience in the paper, the Maynard Smith Replicator Dynamic Smith (1982) (MRD):
@@ 3.2. Learning rate and batch-size
The population-policy equivalence gives an interesting interpretation to the learning rate $\alpha$ commonly used in RL. Under the population perspective, $\alpha$ describes the number of individuals in the population. In Sec. 5, we empirically show that the CL update rule described in Eq. 1 does not typically converge to the optimal action. However, using a small enough learning rate (i.e., a large enough population size) alleviates this issue [^6] As implied by Eq. 14, assuming that, when performed sequentially on randomly sampled individuals instead of one step parallel updates across the entire population, the voter rule still yields the Replicator Dynamic, which we conjecture. We leave a proof of this conjecture for future work..
@@ 4. Methods
We present two types of experiments to validate the findings from the previous section. First, we implement the two RL update rules, CL and MCL in two variants: batched and non-batched. Second, we conduct population-based experiments using $R_{\text{voter}}$ and $R_{\text{wvoter}}$ (VR, WVR) for different population sizes. Moreover, we also numerically simulate the TRD and MRD to show how the above experiments compare with the analytical solutions. It should be noted that MCL is not a competitive bandit algorithm but rather an academic example designed solely to demonstrate the possibility of deriving a biologically plausible RL algorithm equivalent to its SI counterpart, WVR.
@@ 4.2. RL experiments
Moreover, since this update rule can make $\pi$ invalid, i.e., components could become negative or above one (see footnote 8), we clamp $\pi$ between 0 and 1:
@@
These compurtations are carried out for every training step, and there are $S$ steps per seed.
Batched: In these experiments, we implement the batched variants of update rules CL (Eq. 2) and MCL (Eq. 25), henceforth named B-CL and B-MCL. B-CL is a straightforward batching of the CL update rule, averaging over a batch of $B$ samples instead of one sample to update $\pi$. Whereas, with B-MCL, $v^{\pi}$ is no longer a moving average of the rewards but rather the mean of batch rewards. We also need to explicitly clamp the policy between 0 and 1 to ensure it remains valid. B-MCL also uses $B$ samples simultaneously similar to B-CL. Similar to non-batched experiments, we perform $S$ steps per seed.
@@ 5. Results
Finally, from the discussions related to batched RL updates and population experiments, we have empirically validated Proposition 1 and Proposition 2.
@@ 6. Conclusions
With Propositions 1 and 2, we have demonstrated how RD is the underlying connection between Reinforcement Learning and Collective Decision-Making. Further, we have empirically validated this correspondence. This correspondence opens a bridge between these two fields, enabling the flow of ideas, new perspectives, and analogies for both. For example, it can be seen that Cross Learning, Maynard-Cross Learning, and, more generally, Reinforcement Learning takes an individualistic perspective, where information from consecutive action samples/batches is directly accumulated into one centralized agent’s policy. On the other hand, $R_{\text{voters}}$ and $R_{\text{wvoters}}$ take a collectivistic perspective, where every individual implements simple local and decentralized rules, making independent decisions, leading to an emergent collective policy.
@@
Significance for SI. The population-policy equivalence highlights how certain Swarm Intelligence methods are equivalent to single-agent Reinforcement Learning, demonstrating the agency of the entire swarm of non-learning individuals as a single learning entity. Therefore, one could imagine that Multi-Agent Reinforcement Learning would similarly yield equivalent multi-swarm systems, where two or more coexisting swarms would compete/collaborate for resources (i.e., prisoners dilemma, hawk dove, etc). Further, ideas that empower Reinforcement Learning, could be ported to swarm intelligence and swarm robotics. However, extending the demonstrated equivalence to sequential RL would require defining what a state transition means for an entire swarm, which is neither trivial nor straightforward. Contextual bandits could be a possible direction to start thinking about this, where the swarm as a whole would be in a state that determines the options’ qualities/rewards/payoffs. Once the state transitions are defined for the entire swarm, one might benefit from state-couple replicator dynamic Hennes et al. (2009) to derive a similar equivalence for sequential RL.
@@ A.1. Batched RL experiments
Figure 5. Batched RL figures for large batch size and the two other environments where $q^{\pi}_{a}$’s are near zero and near one.
@@ A.2. Population experiments
Figure 6. Population experiments for large population size and the two other environments where $q^{\pi}_{a}$’s are near zero and near one.
@@ A.3. Hyperparametrs
final time ($t_{f}$) $\alpha\times S$ = 100